Required information

[The following information applies to the questions displayed below.]

In 2018, the Westgate Construction Company entered into a contract to construct a road for Santa Clara County for $10,000,000. The road was completed in 2020. Information related to the contract is as follows:

2018 2019 2020
Cost incurred during the year $ 2,100,000 $ 2,450,000 $ 2,695,000
Estimated costs to complete as of year-end 4,900,000 2,450,000 0
Billings during the year 2,200,000 2,350,000 5,450,000
Cash collections during the year 1,900,000 2,300,000 5,800,000


Westgate recognizes revenue over time according to percentage of completion.


rev: 09_15_2017_QC_CS-99734

Required:

1. Calculate the amount of revenue and gross profit (loss) to be recognized in each of the three years. (Do not round intermediate calculations. Loss amounts should be indicated with a minus sign.)
Answer #1

Solution

2018 2019 2020
Revenue $             3,000,000 $        3,500,000 $            3,500,000
Gross profit $                900,000 $        1,050,000 $                805,000

Working

Working 2018 2019 2020
A Contract Price $ 10,000,000 $ 10,000,000 $ 10,000,000
B Cost Incurred to Date $ 2,100,000 $ 4,550,000 $ 7,245,000
C Estimated cost yet to be incurred to complete the contract $ 4,900,000 $ 2,450,000 $ 0
D = B+C Total Cost $ 7,000,000 $ 7,000,000 $ 7,245,000
E = (D/B) x 100 % of Completion 30% 65% 100%
F = A x E Revenue to date $ 3,000,000 $ 6,500,000 $ 10,000,000
G Revenue of Previous year $ 0 $ 3,000,000 $ 6,500,000
H = F - G Net Revenue this year $ 3,000,000 $ 3,500,000 $ 3,500,000
I Cost to date $ 2,100,000 $ 4,550,000 $ 7,245,000
J Cost to date of previous year $ 0 $ 2,100,000 $ 4,550,000
K = I - J Net Cost for the year $ 2,100,000 $ 2,450,000 $ 2,695,000
L = H - K Gross Profits $ 900,000 $ 1,050,000 $ 805,000